# Dynamics on invariant tori emerging through forced symmetry breaking in phase oscillator networks

### v1

@@ 1. Introduction
where $g$ mediates the coupling within populations, while the functions $Y(\theta)=\{Y_{\sigma,k}(\theta)\}$ determine the coupling between populations. We shall assume in this paper that the coupling function $Y_{\sigma,k}$ only depends on within-population phase differences, i.e., variables of the form $\theta_{\tau,j}-\theta_{\tau,k}$. This ensures that (2) is equivariant with respect to a $\mathbb{T}^{m}$-symmetry action, where $\phi=(\phi_{1},\dotsc,\phi_{m})$ acts by
@@
(4) $\begin{split}\dot{\theta}_{\sigma,k}&=\omega+\displaystyle\sum_{j=1}^{n}g(\theta_{\sigma,j}-\theta_{\sigma,k})+Y_{\sigma,k}(\theta)+\delta Z_{\sigma,k}(\theta)\\ &=H_{\sigma,k}(\theta)+\delta Z_{\sigma,k}(\theta)\,,\end{split}$
@@ 2.1. Equivariant dynamics and relative equilibria
(6) $R\phi=(\phi_{1},\ldots,\phi_{1};\phi_{2},\ldots,\phi_{2};\ldots;\phi_{m},\ldots,\phi_{m})\in(\mathbb{T}^{n})^{m}\,.$
@@ 2.3. The linearized dynamics around a relative equilibrium
Proof.
Differentiating $H(\theta+R\phi)=H(\theta)$ to $\theta$ gives $DH(\theta+R\phi)=DH(\theta)$, showing that $DH$ is constant along any group orbit. Setting $\theta=\theta_{0}$ gives $DH(E(\phi))=DH(\theta_{0})$. ∎
@@
Proof.
Differentiating the equivariance identity $H(\theta+R\phi)=H(\theta)$ to $\phi$ at $\phi=0$ gives the result. ∎
@@
    The range of $N$ is normal to the relative equilibrium:
    $\mathbb{R}^{d}={\rm im}\,R\oplus{\rm im}\,N\,.$
@@
    $N$ semi-conjugates $L$ to $DH$ on the relative equilibrium:
    $(DH\circ E)N=NL\,.$
Proof.
Recall that ${\rm im}\,R\subset\ker DH\circ E=\ker DH(\theta_{0})$. The assumption that the relative equilibrium is normally hyperbolic implies that $\ker DH(\theta_{0})={\rm im}\,R$, and also that this kernel can be complemented by the sum of the hyperbolic generalized eigenspaces of $DH(\theta_{0})$. Let $N:\mathbb{R}^{d-m}\to\mathbb{R}^{d}$ be an injective linear map spanning the sum of these generalized eigenspaces (for instance, the matrix of $N$ could have the hyperbolic generalized eigenvectors as its columns.) By construction it then holds that $\mathbb{R}^{d}={\rm im}\,R\oplus{\rm im}\,N$. It also follows that there exists a linear map $L:\mathbb{R}^{d-m}\to\mathbb{R}^{d-m}$ satisfying
@@
The matrix $L$ is called a Floquet matrix and its eigenvalues the Floquet exponents of the invariant torus. We call $L$ hyperbolic if it none of the Floquet exponents lie on the imaginary axis.
@@ 3. Symmetry breaking in coupled phase oscillator networks
We can use the results of the previous section to understand how synchrony patterns in phase oscillators networks (2) perturb under forced symmetry breaking. Specifically, we will focus on $m$ phase oscillator populations of $n$ oscillators each coupled as in [11] where the (unperturbed) dynamics evolve according to
(15) $\begin{split}\dot{\theta}_{\sigma,k}&=\omega+\displaystyle\sum_{j\neq k}\Big{(}g_{2}(\theta_{\sigma,\;j}-\theta_{\sigma,k})-K^{-}G_{4}(\theta_{\sigma-1};\theta_{\sigma,\;j}-\theta_{\sigma,k})\\ &\qquad\qquad\qquad\qquad+K^{+}G_{4}(\theta_{\sigma+1};\theta_{\sigma,\;j}-\theta_{\sigma,k})\Big{)}=H_{\sigma,k}(\theta)\end{split}$
with
$G_{4}(\theta_{\tau};\varphi)=\frac{1}{n^{2}}\sum_{p,q=1}^{n}g_{4}(\theta_{\tau,p}-\theta_{\tau,q}+\varphi),$
where $g_{2}:\mathbb{T}\to\mathbb{R}$ mediates the pairwise interactions within populations and $g_{4}:\mathbb{T}\to\mathbb{R}$ nonpairwise interactions between the populations with strength $K^{-},K^{+}>0$. For concreteness, we set $K:=K^{-}=K^{+}$ and consider the coupling functions
(16a) $g_{2}(\vartheta)$ $=\sin(\vartheta+\alpha_{2})-r_{0}\sin(2(\vartheta+\alpha_{2})),$
@@
where $\alpha_{2},\alpha_{4}\in\mathbb{R}$ are phase shift parameters for the pairwise and nonpairwise interactions respectively. These equations can be motivated by phase reductions of coupled nonlinear oscillators; see, for example, [21, 22]
These equations are equivariant with respect to the phase-shift symmetry action (3). This implies that the synchrony patterns where each population is either synchronized ($\theta_{\sigma}\in\mathrm{S}$) or desynchronized ($\theta_{\sigma}\in\mathrm{D}$),
$\mathrm{X}_{1}\dotsc\mathrm{X}_{m}:=\{\theta=(\theta_{1},\dotsc,\theta_{m})\in\mathbb{T}^{mn}:\theta_{\sigma}\in\mathrm{X}_{\sigma}\in\{\mathrm{S},\mathrm{D}\}\}$
are equilibria relative to the symmetry action of $\mathbb{T}^{m}$. For example, $\mathrm{S}\dotsc\mathrm{S}$ ($m$-times) is the relative equilibrium where all populations are phase synchronized individually (but not necessarily to each other).
For the remainder of the paper, we will focus synchrony patterns in small networks of $m=3$ population of $n=2$ oscillators as considered in [11, 23]. However, the same methods also apply to networks with populations [24] or with more oscillators per population [11].
3.1. Invariant tori as relative equilibria
The dynamics (15) of $m=3$ populations of $n=2$ oscillators define a dynamical system on $\mathbb{T}^{mn}=\mathbb{T}^{6}$ that is equivariant with respect to the continuous $\mathbb{T}^{m}=\mathbb{T}^{3}$ symmetry (3). Write $\mathbb{Z}_{m}:=\mathbb{Z}/m\mathbb{Z}$ for the cyclic group of $m$ elements and $S_{n}$ for the group of permutations of $n$ elements. In addition to the continuous symmetries, the system has two discrete symmetries: The dynamical equations (15) are equivariant with respect to an action of $\mathbb{Z}_{m}$ that acts by permuting the population index $\sigma$ and $(S_{n})^{m}$ that acts by permuting the indices $k$ within populations.
While the continuous symmetries imply that the tori $\mathrm{SSD}$, $\mathrm{SDD}$, $\mathrm{DSS}$, $\mathrm{DSD}$, $\mathrm{SDS}$, $\mathrm{DDS}$ are relative equilibria of (15), the $\mathbb{Z}_{3}$-symmetry allows to restrict attention to two of them: $\mathrm{DSS},\mathrm{SDS}$ are the images of $\mathrm{SSD}$ under the $\mathbb{Z}_{3}$ action and, similarly, there are two symmetric copies of $\mathrm{SDD}$. Thus, to understand the dynamics on these tori it is sufficient to consider $\mathrm{SDD},\mathrm{SSD}$. Both $\mathrm{SDD},\mathrm{SSD}$ are copies of $\mathbb{T}^{3}$ embedded into $\mathbb{T}^{6}$, respectively, through
@@ 3.1. Invariant tori as relative equilibria
and
(18) $E^{\tiny\mathrm{SSD}}(\phi_{1},\phi_{2},\phi_{3})$ $=(\phi_{1},\phi_{1};\phi_{2},\phi_{2};\phi_{3},\phi_{3}+\pi),$
respectively. The vector field on $\mathbf{T}\in\{\mathrm{SDD},\mathrm{SSD}\}$ is constant,
$\Omega^{\mathbf{T}}=(\Omega^{\mathbf{T}}_{1},\Omega^{\mathbf{T}}_{2},\Omega^{\mathbf{T}}_{3})^{\mathsf{T}}.$
Thus, solutions to (15) describe (quasi) periodic motion relative to the $\mathbb{T}^{3}$ symmetry.
While we do not consider the tori SSS and DDD here (as they are not part of the heteroclinic cycles found in [11]), they can be analyzed in the same way.
@@ 3.2. Perturbed tori and dynamics thereon
Our main goal is to understand how the normally hyperbolic invariant tori $\mathrm{SSD},\mathrm{SDD}$ deform under forced symmetry breaking of the continuous symmetry. Specifically, for small $\delta>0$, we consider the perturbed dynamical system
(19) $\dot{\theta}_{\sigma,k}=H_{\sigma,k}(\theta)+\delta Z_{\sigma,k}(\theta),$
and for $\mathbf{T}\in\{\mathrm{SSD},\mathrm{SDD}\}$ seek (first-order) perturbed embeddings
@@
While our approach is more general, we will focus for concreteness on perturbations of the form
where $h$ is a (nonconstant) $2\pi$-periodic function. This corresponds to Kuramoto-type coupling between any two oscillators as in (1), not just within populations. While the equations for $\delta=0$ are $\mathbb{T}^{3}$-equivariant, for $\delta>0$ the system (19) is $\mathbb{T}$-equivariant corresponding to a single phase-shift symmetry as in a single population of Kuramoto-type oscillators (1). Hence, this is an example of forced symmetry breaking. Note that with the perturbation (22) the system (19) is still $\mathbb{Z}_{3}$-equivariant and we can—as for the unperturbed system—focus on how the two tori $\mathrm{SSD},\mathrm{SDD}$ perturb.
To compute the perturbed dynamics and embedding, we solve the homological equations (13). We now state the main results and give details of these computations in the following section.
@@
Consider the phase oscillator network (19) with unperturbed vector field (15), coupling functions (16) for $(\alpha_{2},\alpha_{4})=(\frac{\pi}{2},\pi)$, and perturbation given by (22) with function $h$.
@@
Consider the perturbation problem as above with function
@@
These dynamical equations, together with the approximations of the embeddings of SSD and SDD, will allow to generate starting data to explore solutions of the symmetry broken system numerically. We will do this in Section 5 below.
@@ 4.1. Preliminaries
To compute the first-order approximation of the perturbed tori, we solve the homological equations (13); we first collect the basic ingredients. Written as a matrix, the map $R$ in (6) that determines the phase-shift symmetries of the unperturbed system—and thus directions tangent to the unperturbed torus—is
@@
The normal directions of the torus “split” each oscillator population, that is,
@@
First consider the unperturbed torus $\mathrm{SDD}$. The Floquet matrix that determines the linear dynamics in the normal directions is
@@
Substituting $\mathrm{SDD}$ into (15) with coupling (16) yields the frequency vector $\Omega^{\mathrm{SDD}}=(0,-2,-2)^{\mathsf{T}}$. Note that populations two and three are resonant. Thus, a normal form for the reduced dynamics on the perturbation only depends on $\phi_{1}$ and the phase difference $\phi_{3}-\phi_{2}$; cf. (23) in Proposition 3.1.
For $\mathrm{SSD}$, we obtain in a similar way the Floquet matrix
(31) $L^{\mathrm{SSD}}=\left(\begin{matrix}-4r_{0}+2K&0&0\\ 0&-4r_{0}-2K&0\\ 0&0&-4r_{0}\end{matrix}\right)$
and the frequency vector $\Omega^{\mathrm{SSD}}=(0,0,-2)^{\mathsf{T}}$. Thus the normal form equations for the reduced dynamics will only depend on $\phi_{2}-\phi_{1}$ and $\phi_{3}$.

### v2

@@ Abstract.
Key words and phrases:
Continuous symmetries, relative equilibrium, forced symmetry breaking, invariant torus, numerical continuation.
1991 Mathematics Subject Classification:
Primary: 37C81, 37G40; Secondary: 37D10, 34A12.
Corresponding author: Christian Bick
[…]
1Department of Mathematics, Vrije Universiteit Amsterdam, De Boelelaan 1111, Amsterdam, The Netherlands
2Institute for Advanced Study, Technical University of Munich, Lichtenbergstr 2, 85748 Garching, Germany
3Department of Mathematics, University of Exeter, Exeter EX4 4QF, United Kingdom
4Mathematical Institute, University of Oxford, Oxford OX2 6GG, United Kingdom
@@ 1. Introduction
where $g$ mediates the coupling within populations, while the functions $Y(\theta)=\{Y_{\sigma,k}(\theta)\}$ determine the coupling between populations. We assume in this paper that the coupling function $Y_{\sigma,k}$ only depends on within-population phase differences, i.e., variables of the form $\theta_{\tau,j}-\theta_{\tau,k}$. This ensures that (2) is equivariant with respect to a $\mathbb{T}^{m}$-symmetry action, where $\phi=(\phi_{1},\dotsc,\phi_{m})$ acts by
@@
[…]
@@ 2.1. Equivariant dynamics and relative equilibria
(6) $R\phi=(\phi_{1},\dotsc,\phi_{1};\phi_{2},\dotsc,\phi_{2};\dotsc;\phi_{m},\dotsc,\phi_{m})\in(\mathbb{T}^{n})^{m}\,.$
@@
Note that we use the word “semi-conjugacy” as it is used in [20]: a non-invertible map sending solutions to solutions (in this case an embedding).
@@ 2.3. The linearized dynamics around a relative equilibrium
Proof of Proposition 2.2.
Differentiating $H(\theta+R\phi)=H(\theta)$ with respect to $\theta$ gives $DH(\theta+R\phi)=DH(\theta)$, showing that $DH$ is constant along any group orbit. Setting $\theta=\theta_{0}$ gives $DH(E(\phi))=DH(\theta_{0})$. ∎
@@
Proof of Proposition 2.3.
Differentiating the equivariance identity $H(\theta+R\phi)=H(\theta)$ with respect to $\phi$ at $\phi=0$ gives the result. ∎
@@
    The range of $N$ is normal to the relative equilibrium: R^d = im  R ⊕im  N  .
@@
    NN semi-conjugates $L$ to $DH$ on the relative equilibrium: (DHE) N = N L  .
Proof of Proposition 2.4.
Recall that ${\rm im}\,R\subset\ker DH\circ E=\ker DH(\theta_{0})$. The assumption that the relative equilibrium is normally hyperbolic implies that $\ker DH(\theta_{0})={\rm im}\,R$, and also that this kernel can be complemented by the sum of the hyperbolic generalized eigenspaces of $DH(\theta_{0})$. Let $N:\mathbb{R}^{d-m}\to\mathbb{R}^{d}$ be an injective linear map spanning the sum of these generalized eigenspaces (for instance, the matrix $N$ could have the hyperbolic generalized eigenvectors as its columns). By construction it then holds that $\mathbb{R}^{d}={\rm im}\,R\oplus{\rm im}\,N$. It also follows that there exists a linear map $L:\mathbb{R}^{d-m}\to\mathbb{R}^{d-m}$ satisfying
@@
The injective matrix $N$ in Proposition 2.4 is clearly not unique: we only require that its image is the hyperbolic subspace of $DH(\theta_{0})$. As a result, $L$ is not unique either, although the eigenvalues of $L$ are fixed and equal to the nonzero eigenvalues of $DH(\theta_{0})$. A matrix $L$ satisfying the conclusion of the proposition is called a Floquet matrix for the invariant torus $E(\mathbb{T}^{m})$. Its (unique) eigenvalues are called the Floquet exponents of the invariant torus. The normal hyperbolicity of the torus implies that none of these Floquet exponents lie on the imaginary axis.
@@ 3. Symmetry breaking in coupled phase oscillator networks
We now exploit these results to elucidate the effect of forced symmetry breaking in $m$ coupled populations of $n$ oscillators given by the perturbation problem (4); such phase equations can be motivated by phase reductions [22, 23]. The unperturbed equations for $\delta=0$ are equivariant with respect to the action of $\mathbb{T}^{m}$ given by (3), which corresponds to a phase-shift symmetry in each population. This implies that the synchrony patterns where each population is either synchronized ($\theta_{\sigma}\in\mathrm{S}$) or desynchronized ($\theta_{\sigma}\in\mathrm{D}$),
(15) $\mathrm{X}_{1}\dotsc\mathrm{X}_{m}:=\{\theta=(\theta_{1},\dotsc,\theta_{m})\in\mathbb{T}^{mn}:\theta_{\sigma}\in\mathrm{X}_{\sigma}\in\{\mathrm{S},\mathrm{D}\}\}$
are equilibria relative to the symmetry action of $\mathbb{T}^{m}$. For example, $\mathrm{S}\dotsc\mathrm{S}$ ($m$-times) is the relative equilibrium where all populations are phase synchronized individually (but not necessarily to each other). Perturbing the system, $\delta>0$, breaks the $\mathbb{T}^{m}$ symmetry to a $\mathbb{T}$ symmetry where $\alpha\in\mathbb{T}$ acts by a common phase shift of all phases, $\theta_{\sigma,k}\mapsto\theta_{\sigma,k}+\alpha$. We can now compute how invariant tori (15) and the dynamics deform as the symmetry is broken.
While the results apply more generally to systems of the form (4), we focus on the case of $m=3$ population of $n=2$ oscillators as considered in [11, 24]. Specifically, the oscillators in population $\sigma$ evolve according to
(16a) $\dot{\theta}_{\sigma,1}$ $=H_{\sigma,1}(\theta)+\delta Z_{\sigma,1}(\theta)=\omega+g(\theta_{\sigma,2}-\theta_{\sigma,1})+Y_{\sigma,1}(\theta)+\delta Z_{\sigma,1}(\theta)$
(16b) $\dot{\theta}_{\sigma,2}$ $=H_{\sigma,2}(\theta)+\delta Z_{\sigma,2}(\theta)=\omega+g(\theta_{\sigma,1}-\theta_{\sigma,2})+Y_{\sigma,2}(\theta)+\delta Z_{\sigma,2}(\theta)$
where the coupling between populations with strength $K\geq 0$ is determined by
(17) $\begin{split}Y_{\sigma,k}(\theta)&=-\frac{K}{4}\big(g_{4}(\theta_{\sigma-1,1}-\theta_{\sigma-1,2}+\theta_{\sigma,3-k}-\theta_{\sigma,k})\\ &\qquad\qquad+g_{4}(\theta_{\sigma-1,2}-\theta_{\sigma-1,1}+\theta_{\sigma,3-k}-\theta_{\sigma,k})\big)\\ &\ \quad+\frac{K}{4}\big(g_{4}(\theta_{\sigma+1,1}-\theta_{\sigma+1,2}+\theta_{\sigma,3-k}-\theta_{\sigma,k})\\ &\qquad\qquad+g_{4}(\theta_{\sigma+1,2}-\theta_{\sigma+1,1}+\theta_{\sigma,3-k}-\theta_{\sigma,k})\big)\end{split}$
for $k=1,2$ and $\sigma=1,2,3$ taken modulo $3$. The $2\pi$-periodic coupling functions
(18a) $g(\vartheta)=g_{2}(\vartheta)$ $=\sin(\vartheta+\alpha_{2})-r_{0}\sin(2(\vartheta+\alpha_{2})),$
@@
with parameters $\alpha_{2},\alpha_{4}\in\mathbb{R}$ mediate the pairwise interactions within populations and nonpairwise interactions between populations, respectively. For $\delta=0$, the system (16) is $\mathbb{T}^{3}$ equivariant.
For $\delta>0$ the network dynamics (16) is now subject to forced symmetry breaking. We focus on a symmetry breaking perturbation of the form
where $h$ is a (nonconstant) $2\pi$-periodic function. This corresponds to pairwise Kuramoto-type coupling between any two oscillators as in (1) and breaks the phase shift symmetries induced by the nonpairwise coupling between populations: The $\mathbb{T}^{3}$ symmetry is broken to a single $\mathbb{T}$ symmetry. Together, Equations (16)–(19) define the system we investigate.
3.1. Invariant tori as relative equilibria for $\delta=0$
Because of the phase-shift symmetry, we may assume $\omega=-1$ without loss of generality [^1] While this value is different from the more common choice of $\omega=0$, it is convenient for plotting as synchronized solutions appear stationary.. In addition to the continuous symmetries, the system also admits discrete symmetries: Write $\mathbb{Z}_{m}:=\mathbb{Z}/m\mathbb{Z}$ for the cyclic group of $m$ elements. The dynamical equations (16) are equivariant with respect to an action of $\mathbb{Z}_{3}$ that acts by permuting the population index $\sigma$ and $(\mathbb{Z}_{2})^{3}$ that acts by permuting the indices $k$ within populations.
The unperturbed dynamics for $\delta=0$ have eight invariant tori, $\mathrm{SSS},\mathrm{DDD}$ as well as $\mathrm{SSD},\mathrm{SDD},\mathrm{DSS},\mathrm{DSD},\mathrm{SDS},\mathrm{DDS}$. We focus on the latter as they are part of the heteroclinic cycles found in [11]. For concreteness, we set $(\alpha_{2},\alpha_{4})=(\frac{\pi}{2},\pi)$ in the parameter range where where heteroclinic cycles involving these tori may exist. We can further restrict our attention to $\mathrm{SDD},\mathrm{SSD}$ as the remaining tori are their images under $\mathbb{Z}_{3}$ action.
The torus $\mathrm{SDD}$ is a copy of $\mathbb{T}^{3}$ embedded into $\mathbb{T}^{6}$ through the map
@@ 3.1. Invariant tori as relative equilibria for $\delta=0$
Since $g_{2}(0)=1$, $g_{2}(\pi)=-1$ and $g_{4}(0)=g_{4}(\pi)=0$, the constant vector field on $\mathrm{SDD}$ is given by $\Omega^{\mathrm{SDD}}=(0,-2,-2)^{\mathsf{T}}$. Transverse stability of the invariant torus is governed by the eigenvalues corresponding to the directions that break $\mathrm{S}$, $\mathrm{D}$ in each direction (e.g., the eigenvector (1,1;0,0;0,0)(-1,1;0,0;0,0) for the first population); cf. [11]. Taking derivatives these evaluate to $\lambda_{1}^{\mathrm{SDD}}=-4r_{0}$, $\lambda_{2}^{\mathrm{SDD}}=-4r_{0}+2K$, and $\lambda_{3}^{\mathrm{SDD}}=-4r_{0}-2K$.
Similarly, the torus $\mathrm{SSD}$ can be parameterized by the embedding
(21) $E^{\tiny\mathrm{SSD}}(\phi_{1},\phi_{2},\phi_{3})$ $=(\phi_{1},\phi_{1};\phi_{2},\phi_{2};\phi_{3},\phi_{3}+\pi).$
The constant dynamics on the torus are determined by the frequency vector $\Omega^{\mathrm{SSD}}=(0,0,-2)^{\mathsf{T}}$. Transverse stability is governed by the eigenvalues $\lambda_{1}^{\mathrm{SSD}}=-4r_{0}+2K$, $\lambda_{2}^{\mathrm{SSD}}=-4r_{0}-2K$, and $\lambda_{3}^{\mathrm{SSD}}=-4r_{0}$.
@@ 3.2. Perturbed tori and dynamics thereon
Our main goal is to understand how the normally hyperbolic invariant tori $\mathrm{SSD},\mathrm{SDD}$ for (16) deform under forced symmetry breaking for $\delta>0$. Thus, for $\mathbf{T}\in\{\mathrm{SSD},\mathrm{SDD}\}$ we seek (first-order) perturbed embeddings
@@
To compute the perturbed dynamics and embedding for $\delta>0$, we solve the homological equations (13). We now state the main results and give details of these computations in the following section.
@@
Consider the phase oscillator network (16) of $m=3$ populations with $n=2$ oscillators each for $(\alpha_{2},\alpha_{4})=(\frac{\pi}{2},\pi)$ subject to the perturbation (19).
@@
Consider the perturbation problem (16) as above with function
@@
These dynamical equations, together with the approximations of the embeddings of SSD and SDD, will allow us to generate starting data to explore solutions of the symmetry broken system numerically. We will do this in Section 5 below.
@@ 4.1. Preliminaries
To compute the first-order approximation of the perturbed tori, we solve the homological equations (13). We first collect the basic ingredients from Section 3.1. Written as a matrix, the map $R$ in (6) that determines the phase-shift symmetries of the unperturbed system—and thus directions tangent to the unperturbed torus—is
@@
Because the Jacobian matrices of (16) (for $\delta=0$) at SSD and SDD are symmetric matrices, the images of $R$ and $N$ are orthogonal. An example of a matrix $N$ of which the image is orthogonal to that of $R$ is given by
@@
The rows of $N$ represent a “splitting” of each oscillator population. The Floquet matrix that determines the linear dynamics in the normal directions is now computed by solving the linear equation $(DH)N=NL$, where $DH$ is the Jabobian of the unperturbed ODE given in (16), evaluated at SSD or SDD.
The entries of the Floquet matrices relate directly to the transverse stability of the relative equilibria (given in Section 3.1). Thus, for the unperturbed torus $\mathrm{SDD}$, we compute
@@
Since the unperturbed dynamics are given by the frequency vector $\Omega^{\mathrm{SDD}}=(0,-2,-2)^{\mathsf{T}}$, populations two and three are resonant. Thus, a normal form for the reduced dynamics on the perturbation only depends on $\phi_{1}$ and the phase difference $\phi_{3}-\phi_{2}$; cf. (24) in Proposition 3.1. Similarly, for $\mathrm{SSD}$ we obtain in a similar way the Floquet matrix
(32) $L^{\mathrm{SSD}}=\left(\begin{matrix}-4r_{0}+2K&0&0\\ 0&-4r_{0}-2K&0\\ 0&0&-4r_{0}\end{matrix}.\right)$
With the frequency vector $\Omega^{\mathrm{SSD}}=(0,0,-2)^{\mathsf{T}}$ the normal form equations for the reduced dynamics will only depend on $\phi_{2}-\phi_{1}$ and $\phi_{3}$.